$SPAD/src/input tutchap1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool tutchap1.output
)set message test on
)set message auto off
)clear all
--S 1 of 19
1+1
--R
--R
--R   (1)  2
--R                                                        Type: PositiveInteger
--E 1
--S 2 of 19
123^45
--R
--R
--R   (2)
--R  1111040818513195628591079058717645191855915321226802182362907319986611100124_
--R   2743283966127048043
--R                                                        Type: PositiveInteger
--E 2
--S 3 of 19
2^(3+4)
--R
--R
--R   (3)  128
--R                                                        Type: PositiveInteger
--E 3
--S 4 of 19
4/3
--R
--R
--R        4
--R   (4)  -
--R        3
--R                                                      Type: Fraction(Integer)

--E 4
--S 5 of 19
2/2
--R
--R
--R   (5)  1
--R                                                      Type: Fraction(Integer)
--E 5
--S 6 of 19
a := 2
--R
--R
--R   (6)  2
--R                                                        Type: PositiveInteger
--E 6
--S 7 of 19
b := a
--R
--R
--R   (7)  2
--R                                                        Type: PositiveInteger
--E 7
--S 8 of 19
a := 3
--R
--R
--R   (8)  3
--R                                                        Type: PositiveInteger
--E 8
--S 9 of 19
b
--R
--R
--R   (9)  2
--R                                                        Type: PositiveInteger
--E 9
--S 10 of 19
i : Integer
--R
--R                                                                   Type: Void
--E 10
--S 11 of 19
i := 2/3

--R
--R
--RDaly Bug
--R   Cannot convert right-hand side of assignment
--R   2
--R   -
--R   3
--R
--R      to an object of the type Integer of the left-hand side.
--E 11
--S 12 of 19
i := 4/2
--R
--R
--R   (11)  2
--R                                                                Type: Integer
--E 12
--S 13 of 19
c : PositiveInteger := 3
--R
--R
--R   (12)  3
--R                                                        Type: PositiveInteger
--E 13
--S 14 of 19
(j, k, l) : Integer
--R
--R                                                                   Type: Void
--E 14
--S 15 of 19
)display names
--R
--R
--RNames of User-Defined Objects in the Workspace:
--R
--R%    a    b    c    i    j    k    l
--R
--RNames of System-Defined Objects in the Workspace:
--R
--R%e                %i                %infinity         %minusInfinity
--R%pi               %plusInfinity     SF
--E 15
)clear properties c i
--S 16 of 19

%%(1)
--R
--R
--R   (14)  2
--R                                                        Type: PositiveInteger
--E 16
--S 17 of 19
1 + %%(-3)
--R
--R
--R   (15)  4
--R                                                        Type: PositiveInteger
--E 17
--S 18 of 19
1 _
  +_
  2
--R
--R
--R   (16)  3
--R                                                        Type: PositiveInteger
--E 18
--S 19 of 19
7 * 8 -- In the next chapter we shall move beyond elementary arithmetic.
--R
--R
--R   (17)  56
--R                                                        Type: PositiveInteger
--E 19
)spool
)lisp (bye)

———-

References^{[1] nothing}

7